University Florists makes bouquets from a variety of materials. The Daily Special Bouquet is priced at $20. The florist assembles this bouquet each day from a variety of low cost flowers he buys from his flower supplier. The actual cost of flowers ranges uniformly from $2 to $7, with all intermediate values being equally likely. The florist (who studied management science many years ago) knows that the time to assemble a bouquet is normally distributed with a mean time of 5 minutes and standard deviation of 1 minute. This will be the time required for all of the Daily Special Bouquets for that day. The florist values his labor time at $10 per hour. Sales are normally distributed with a mean of 10 bouquets per day with a standard deviation of 1 bouquet. What formulas should go in cells B8:B14 to simulate daily profits for the store? You just have to write the formulas. No solution of the question is needed.

A

B

C

D

E

F

1

Initial Conditions:

2

Sales Price

$20

Constant

3

Number of Bouquets: Average

10

Std dev

1

Normal

4

Cost of Flowers Lo

$2

Hi

$7

Cont. Uniform

5

Labor Time Average

5

Std dev

1

Normal

6

Labor Cost

$10

Constant

7

8

Cost of Flowers

$4.41

9

Labor Time (Min)

2.55

10

Labor Cost

$0.43

11

Total Cost

4.83

12

Profit/Bouquet

$15.17

13

# Bouquets Sold

12.20

14

Total Profit

$184.99
Answer #1

ANS:

Cell Formula

B8 =PsiUniform($C$4, $E$4)

B9 =PsiNormal($C$5, $E$5)

B10 =B9/60*C6

B11 =B8+B10

B12 =C2-B11

B13 =PsiNormal($C$3, $E$3)

B14 =B12*B13+PsiOutput()